OpenStudy (anonymous):

Let f(x) = 4x – 32 and g(x) = 4x. Part 1 What is (f * g)(x) ? Part 2 Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

\[=f(x)g(x)=(4x-32)(4x)=16x^2-128x=16(x^2-8x)\]

OpenStudy (anonymous):

I'm assuming you wanted the product of the functions and not a composite function\[(f o g)(x)\]

OpenStudy (anonymous):

The composite function would be \[f(g(x))=4(4x)-32=16x-32\]